Translate the verbal sentence into an equation or inequality. Then determine if the given number is a solution:

4 more than the quotient of 2 and 3, times a number k is 20. when k = 24

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



The quotient of 2 and 3:

4 more than ... :

...times a number :

...is 20:

Then substitute 24 in place of , do the arithmetic, and see if you have a true statement or not. If true, then 24 is a solution, otherwise not.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by stanbon(75887)

4 more than the quotient of 2 and 3, times a number k is 20. then k = 24
-------
(2/3)k + 4 = 20
-------
(2/3)k = 16
-------
k = 3*8
-------
k = 24
====================
Cheers,
Stan H.
=========================